Mary’s grandmother’s age is 11 years less than 9 times mary’s age. The sum of their ages is 69. Find each of their ages.
Answer by ikleyn(52784)

.

Let Mary's age will be x years.


Then the grandmother's age is  (9x - 11) years.


According to the condition,


    x + (9x-11) = 69.


Simplify and solve for x


    10x = 69 + 11

    10x = 80

      x = 80/10 = 8.


ANSWER.  Mary is 8 years old;  her grandmother is  69-8 = 61 years old.


CHECK.   8*9 - 11 = 72 - 11 = 61 years.  ! Correct !

Solved.